3. Assume x/x .10 and ż/z .02, and suppose that x(0) 2 and z(0) 1.Calculate thenumerical values of y(t) fort1, t 2, and t 10 for the following cases: (a) y - xz (b) y z/x (c) y = xßZ1-β, where β = 1 /3
Answer #1

We have dr/dt = 0.1 or dr 0.1dt , and taking integral on both sides, we have int rac{1}{x}mathrm{d} x = 0.1 int mathrm{d} t or ln(x) = 0.1 t + c_1 or x = e^{0.1 t + c_1} or x = e^{c_1}e^{0.1 t} or x = C_1e^{0.1 t} . Now, for t=0, we have x=2 or 2 = C_1e^{0.1 *0} or C_1 = 2 , and hence, x = 2e^{0.1 t} .

Simlarly, we have d:/dt 0.02 or 0.02ct , and taking integral on both sides, we have int rac{1}{z}mathrm{d} z = 0.02 int mathrm{d} t or In (:) = 0.02t+c2 or 0.02t+e2 or z = e^{c_2}e^{0.02 t} or z = C_2e^{0.02 t} . Now, for t=0, we have z=1 or 1= C_2e^{0.02 *0} or C_2 = 1 , and hence, z = e^{0.02 t} .

(a) For y = xz , we have y = (2e^{0.1 t})(e^{0.02 t}) or у 2e0.1f+0.02 t or y = 2e^{0.12 t} . Hence,

  • y(0) = 2e^{0.12 *0} or y(0) = 2 .
  • y(1) = 2e0.12.1 or y( 1 ) 2e0.12-2.255 .
  • y(2) = 2e0·13.2 or y(2) = 2E024 = 2.5425 .
  • 10) 2e0.12.1 or y(10) = 2e.-_ 6.6402 .

(b) For y = x/z , we have y = (2e^{0.1 t})/(e^{0.02 t}) or y = 2e^{0.1 t - 0.02 t} or y = 2e^{0.08 t} . Hence,

  • y(0) = 2e008.。 or y(0) = 2 .
  • y(1) = 2e^{0.08 *1} or y(1)=2e008 = 2.1666 .
  • y(2) = 2e^{0.08 *2} or y(2) = 2e^{0.16} = 2.347 .
  • y(10) = 2e^{0.08 *10} or y(10)= 2e®8-4.45 11 .

(c) For y = x^eta z^{1-eta} or y = x^{1/3} z^{1-1/3} or 1/3.2/3 , we have y = (2e^{0.1 t})^{1/3}(e^{0.02 t})^{2/3} or y = (2e^{0.1 t/3})(e^{0.04 t/3}) or 0.1/3+0.04/3 or y = 2e^{0.14 t/3} . Hence,

  • y(0) 2e0.14.0/3 or y(0) = 2 .
  • 110.14 13 or y( 1 ) = 2e0.14/3 = 2.0955 .
  • 2)2e0.14.2/s or y(2) = 2e028/3-2, 1957 .
  • 10) 2e0.1 4+10/3 or (10) 2e1.4-3.189:3 .